# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v2

@@ 1.2 Related Work
We now review the bodies of literature closest to ours. We refer interested readers to Bishop and Nasrabadi (2006) and Bishop and Bishop (2023) for a full horizon of classical ML and modern deep learning as we will not cover concepts such as backpropagation and iterative methods to optimize non-convex neural networks, e.g., stochastic gradient descent (Amari 1993).
@@
Convex optimization for machine learning (ML) training is desirable as a convex training program guarantees convergence to the minimal loss. In recent years, Pilanci and Ergen (2020) have shown that training a feedforward shallow neural network (SNN), i.e., a one-hidden layer NN, with $\operatorname{ReLU}$ activation functions is equivalent to a convex training procedure. They also extended their results to convolutional neural networks (CNNs). More importantly, they fundamentally linked the $\operatorname{ReLU}$-SNN training problem to a regularized convex program. Following this work, Mishkin et al. (2022) have obtained an efficient convex training procedure of SNNs with graphics processing unit (GPU) acceleration, and Kuelbs et al. (2024) have designed an adversarial training procedure by controlling the worst-case output of the neural network in a neighbourhood around each training samples. In this work, we generalize the regularization procedure of feedforward SNNs in regression tasks through the lens of distributionally robust optimization with the Wasserstein metric. The metric is introduced in Section 2.1.
@@
Distributionally robust optimization (DRO), in the ML setting, aims to minimize the expected training loss of an ML model over the most adverse data distribution within a distance of the training set, i.e., the distribution inside a specified ambiguity set, constructed from the available training data, that generates the biggest expected loss. As demonstrated by Kuhn et al. (2019), DRO minimizes the worst-case optimal risk associated with the training of the model. Different authors have worked on distributionally robust neural networks. Bai et al. (2023a) propose to train deep neural networks using the first-order approximation of the Wasserstein DRO loss function in the backpropagation but leave the numerical evaluation for future work. Wasserstein DRO uses the definition of the Wasserstein distance to construct the ambiguity set. We cover it in Section 2.1. The authors of Levy et al. (2020) have developed a DRO gradient estimator that integrates into neural networks’ non-convex optimization pipeline similarly to the widely used stochastic gradient descent (SGD) optimizer. In Sagawa et al. (2020), the authors propose a method to train deep NNs in a DRO fashion by defining the ambiguity set with groups of the training data and increasing regularization. These methods integrate distributional robustness in the training procedure of deep NNs but do not directly tackle the Wasserstein DRO problem, or an equivalent reformulation. They also all suffer from the lack of global optimality guarantees making them harder to deploy in large-scale critical applications. Our proposition, while only applying to shallow networks, works with the tractable reformulation of the Wasserstein DRO problem proposed by Mohajerin Esfahani and Kuhn (2018) and converges to a global optimum thanks to convexity.
@@
The training approach for physics-informed neural networks (PINNs) primarily involves a weighted multi-objective function designed to narrow the gap between labels and model predictions while penalizing physical constraint violations. This is achieved by simultaneously minimizing the residuals of a partial differential equation (PDE) and its boundary conditions (Cuomo et al. 2022). Suppose all NN training weights are to be collected in $\bm{\Theta}$, a typical PINN training program takes the form:
@@
where $\mathcal{L_{F}}$ and $\mathcal{L_{B}}$ are the functions enforcing the differential equations and boundary conditions, respectively, of the governing physical system, $\mathcal{L}$ is the regression objective between labeled data and model predictions, and $\hat{\mathbf{Y}}=\Phi(\mathbf{X},\bm{\Theta})$ (Cuomo et al. 2022). The scalar weights $\lambda_{\mathcal{F}}$, $\lambda_{\mathcal{B}}$, and $\lambda_{\text{data}}$ must be tuned empirically. We note the lack of hard constraints. The loss terms related to the PDE are derived through the automatic differentiation of the neural network’s outputs (Baydin et al. 2017). This complex loss function is then iteratively optimized through backpropagation within the network. While authors of Cuomo et al. (2022) note that boundary conditions can be strictly imposed during training by forcing a part of the network to satisfy these conditions, see Zhu et al. (2021), they are typically integrated as penalty terms within the loss function. By using the augmented Lagrangian in the objective function and increasing the penalty terms related to constraint violations, it is possible to empirically converge to a PINN which imposes hard inequality constraints during training (Lu et al. 2021). Physics-informed neural networks are characterized by their extensive modelling capabilities, facilitated by automatic differentiation, that allows for the approximation of continuous PDEs across various neural network architectures, e.g., CNNs (Gao et al. 2021), long short-term memory networks (Zhang et al. 2020), Bayesian NNs (Yang et al. 2021), and generative adversarial networks (Yang et al. 2020). Although limited to convex constraints, in this work, we demonstrate that our training methodology can enforce hard constraints on the shallow network’s output during training without complexifying its architecture or objective function. These constraints may include, but are not limited to, boundary constraints, ramping constraints, and constraints derived from discrete-time dynamical systems.
@@
As stated previously, prediction algorithms used in large-scale VPP applications must respect some inherent constraints. They need to be computationally scalable, they must be able to model nonlinearities, and the sensibility of the application requires trustworthiness in their predictions, e.g., robustness, interpretability, and guarantees. Regarding the forecast of non-residential buildings’ energy consumption, because grid operators do not generally have access to information on building activity, occupancy, and control, the system is approached as a blackbox and data-driven approaches tend to be the default option (Amasyali and El-Gohary 2018). Authors have used support vector regression (SVR) (Dong et al. 2005) and other hybrid methods leveraging both classification and linear regression to satisfy the aforementioned constraints (Amasyali and El-Gohary 2018). These methods are linear by parts and can approximate nonlinear distributions adequately. Statistical methods such as ARIMA (Newsham and Birt 2010) and Gaussian processes have also proved to be good contenders as they even offer empirical uncertainty bounds (Weng and Rajagopal 2015). Many authors have proposed deep-learning methods, yet these propositions do not have much industrial acceptability as they fail to meet the industrial constraints and bloat the requirements for a safe ML pipeline. In this work, we show that our proposition is interesting for virtual power plant applications where the forecast of many distributed assets is used for critical decision-making as we satisfy all industrial constraints better than other comparable models.
@@
Distributionally robust learning, in general, aims at protecting machine learning models from out-of-distribution outliers in the training set. These outliers tend to be adversarial to the prediction quality of ML models in validation and testing sets and being able to ignore them should increase models’ out-of-sample performance. As such, to correctly benchmark distributionally robust machine learning models, one must be able to generate outliers that are indeed out-of-distribution with respect to the original distribution and be able to do it regardless of the form of the original distribution. In the literature, we found that methodologies to achieve these two objectives in regression tasks are limited. For example, Chen and Paschalidis (2018) create a training set by randomly drawing samples from two distributions, the inlying and outlying one, under probability $p\in(0,1)$ and $1-p$, respectively. They used Gaussian distributions for both. In Qi et al. (2022), authors task their model with the forecast of a real-world dataset. To control the adversity of the setting, they discard the real labels and create a synthetic version of them by using the prediction of a linear regression trained on the original dataset. They then tune an additive noise over the labels to simulate different levels of outlier corruption. In Liu et al. (2022), they notably propose generating corrupted synthetic data by integrating attack mechanisms into the generation to bias their dataset, i.e., selection bias and anti-causal bias. While their methodology is interesting, they must rely on multiple Gaussian distributions to design their attacks as intended. By leveraging the sliced-Wasserstein distance (Bonnotte 2013), a tractable approximation of the Wasserstein distance, our methodology generates outliers, does not rely on any assumptions of the original distribution, making it quite flexible, and ensures that outliers do not create tangible patterns.
@@ 1.3 Contributions
We propose a new methodology to benchmark distributionally robust (and regularized) machine learning models in a controlled synthetic environment. We design a method to introduce real out-of-sample outliers under the sliced-Wasserstein distance (Bonnotte 2013), a computationally tractable approximation of the Wasserstein distance. We combine this method with noise to corrupt machine learning datasets generated with standard optimization benchmark functions typically reserved for non-convex solver benchmarking. We show that our WaDiRo-SCNN training performs well empirically in highly corrupted environments: it is robust and conservative while being able to model non-convex functions. Finally, we test our model on a real-world VPP application, viz., the hourly energy consumption prediction of non-residential buildings from Montréal, Canada. For this application, we enforce the physical constraint that energy consumption is non-negative. We observe that this constraining greatly reduces the number of constraint violations during testing while having no significant precision drop with the original non-physics constrained WaDiRo-SCNN.
@@
The rest of the paper is organized as follows. In Section 2, we introduce Wasserstein distributionally robust optimization and its tractable formulation for convex problems (Section 2.1) as well as shallow convex neural networks (Section 2.2). In Section 3 we present the main results of this paper, viz., WaDiRo-SCNNs. In Sections 4, 5, and 6 we address the theoretical out-of-sample performance of our proposition, physics-constrained SCNNs, and SCNN post-training verification, respectively. Finally, in Section 7 we present a methodology to generate outliers and in Section 8, we present all numerical experiments.
@@ 2.1 Distributionally Robust Optimization
$\inf_{\bm{\beta}}\sup_{\mathbb{Q}\in\Omega}\ \mathbb{E}^{\mathbb{Q}}[h_{\bm{\beta}}(\mathbf{z})],$ (DRO)
where $\mathbf{z}\in\mathcal{Z}\subseteq\mathbb{R}^{d},d\in\mathbb{N}$, is a random sample, $\mathcal{Z}$ is the set of all possible values of $\mathbf{z}$, $\mathbb{Q}$ is a distribution from the ambiguity set $\Omega$, $\bm{\beta}\in\mathbb{R}^{p}$ is the $p$-dimension decision vector, and $h_{\bm{\beta}}(\mathbf{z}):\mathcal{Z}\times\mathbb{R}^{p}\to\mathbb{R}$ is the objective function when applying $\bm{\beta}$ on a sample (Chen and Paschalidis 2020, Mohajerin Esfahani and Kuhn 2018).
We are interested in a data-driven branch of DRO which defines the ambiguity set as a Wasserstein ball, a ball which contains all distributions within a Wasserstein distance from a reference distribution (Yue et al. 2022), centered on an empirical distribution containing $N$ samples: ^$\hat{\mathbb{P}}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{z}_{i}}(\mathbf{z})$, where $\delta_{\mathbf{z}_{i}}(\cdot)$ is a Dirac delta function assigning a probability mass of one at each known sample $\mathbf{z}_{i}$. Note that as the number of samples goes to infinity the empirical distribution converges to the true distribution $\mathbb{P}$ (van der Vaart 2000, Chapter 19.2). The Wasserstein distance, also called earth mover’s distance and closely related to the optimal transport plan (Kuhn et al. 2019), can be pictured as the minimal effort that it would take to displace a given pile of a weighted resource, e.g., dirt, to shape a second specific pile. The effort of a trip is given by the distance traveled and the weight of the carried resource. Specifically, it provides a sense of similarity between the two distributions. Formally, the probabilistic definition of the order-$t$ Wasserstein distance between probability measures $\mathbb{U}$ and $\mathbb{V}$ on $\mathcal{Z}$ is:
$W_{\|\cdot\|,t}(\mathbb{U},\mathbb{V})=\inf_{\begin{subarray}{c}U\sim\mathbb{U},\\ V\sim\mathbb{V}\end{subarray}}\left(\mathbb{E}[\|U-V\|^{t}]\right)^{\frac{1}{t}},\quad t\geq 1,$
where the infimum is taken over all possible couplings of random vectors $U$ and $V$ which are marginally distributed as $\mathbb{U}$ and $\mathbb{V}$, respectively (Panaretos and Zemel 2019). As such, the Wasserstein ball defines a space containing all possible distributions close to the distribution in its center. We refer interested readers to the open-access book of Panaretos and Zemel (2020) which covers the Wasserstein distance thoroughly while offering interesting statistical insights.
@@
$\inf_{\beta}\sup_{\mathbb{Q}\in\Omega}$ $\mathbb{E}^{\mathbb{Q}}[h_{\beta}(\mathbf{z})]$ (WDRO)
$\text{with}\ \Omega=$$\{\mathbb{Q}\in\mathcal{P}(\mathcal{Z}):W_{\|\cdot\|,t}(\mathbb{Q},\hat{\mathbb{P}}_{N})\leq\epsilon\},$
where $\Omega$ is the order-$t$ Wasserstein ball of radius $\epsilon>0$ centered on ^$\hat{\mathbb{P}}_{N}$, and $\mathcal{P}(\mathcal{Z})$ is the space of all probability distributions with support on $\mathcal{Z}$. In a way, $\epsilon$ can be interpreted as a hyperparameter that controls our conservatism towards ^$\hat{\mathbb{P}}_{N}$ since a greater radius admits distributions farther from ^$\hat{\mathbb{P}}_{N}$ into the ambiguity set. This problem is not tractable (Mohajerin Esfahani and Kuhn 2018).
@@
$\inf_{\bm{\beta}}\sup_{\mathbb{Q}\in\Omega}\ \mathbb{E}^{\mathbb{Q}}[h_{\bm{\beta}}(\mathbf{z})]=$ $\inf_{\bm{\beta}}\kappa\epsilon+\frac{1}{N}\sum_{i=1}^{N}h_{\bm{\beta}}(\mathbf{z}_{i})$ (TWDRO)
[…]
where $\|\cdot\|_{*}$ is the dual norm $\|\bm{\theta}\|_{*}\equiv\sup_{\|\mathbf{z}\|\leq 1}\bm{\theta}^{\top}\mathbf{z}$ and $h_{\bm{\beta}}^{*}(\bm{\theta})$ is the convex conjugate of $h_{\bm{\beta}}(\mathbf{z})$, $h_{\bm{\beta}}^{*}(\bm{\theta})=\sup_{\mathbf{z}}\{\bm{\theta}^{\top}\mathbf{z}-h_{\bm{\beta}}(\mathbf{z})\}$ (Boyd and Vandenberghe 2004).
Training machine learning models with (TWDRO) has been proved to be a generalization of regularization (Shafieezadeh-Abadeh et al. 2019). By training over the worst-case expected distribution in a Wasserstein ball centered on the training dataset, we can expect to avoid overfitting on the empirical distribution and yielding increased robustness against outliers, noise, and some adversarial attacks.

### v3

@@ 1.2 Related Work
We now review the strands of literature closest to ours. We refer interested readers to Bishop and Nasrabadi (2006) and Bishop and Bishop (2023) for a full horizon of classical ML and modern deep learning as we will not cover concepts such as backpropagation and iterative methods to optimize non-convex neural networks, e.g., SGD (Amari 1993).
@@
Convex optimization for machine learning (ML) training procedures is desirable as it guarantees convergence to the minimal loss. In recent years, Pilanci and Ergen (2020) have shown that training a feedforward shallow neural network (SNN), i.e., a one-hidden-layer NN, with $\operatorname{ReLU}$ activation functions is equivalent to a convex training procedure. They also extended their results to convolutional neural networks (CNNs). More importantly, they fundamentally linked the $\operatorname{ReLU}$-SNN training problem to a regularized convex program. Following this work, Mishkin et al. (2022) have obtained an efficient convex training procedure of SNNs with graphics processing unit (GPU) acceleration, and Kuelbs et al. (2024) have designed an adversarial training procedure by controlling the worst-case output of the neural network in a neighbourhood around each training sample. In this work, we generalize the regularization procedure from the convex training of Pilanci and Ergen (2020) through the lens of distributionally robust optimization with the Wasserstein metric. The metric is introduced in Section 2.1.
@@
Distributionally robust optimization (DRO), in the ML setting, aims to minimize the expected training loss of an ML model over the most adverse data distribution within a distance of the training set, i.e., the distribution inside a specified ambiguity set, constructed from the available training data, that generates the biggest expected loss. As demonstrated by Kuhn et al. (2019), DRO minimizes the worst-case optimal risk associated with the training of the model. Different authors have worked on distributionally robust neural networks. Bai et al. (2023a) propose to train deep neural networks using the first-order approximation of the Wasserstein DRO loss function in the backpropagation but leave the numerical evaluation for future work. Wasserstein DRO uses the definition of the Wasserstein distance to construct the ambiguity set. We cover it in Section 2.1. The authors of Levy et al. (2020) have developed a DRO gradient estimator that integrates into neural networks’ non-convex optimization pipeline similarly to the widely used SGD optimizer. In Sagawa et al. (2020), the authors propose a method to train deep NNs in a DRO fashion by defining the ambiguity set with groups of the training data and increasing regularization. These methods integrate distributional robustness in the training procedure of deep NNs but do not directly tackle the Wasserstein DRO problem, or an equivalent reformulation. They also all suffer from the lack of global optimality guarantees, making them harder to deploy in large-scale critical applications. Our proposition, while only applying to shallow networks, builds on the tractable reformulation of the Wasserstein DRO problem proposed by Mohajerin Esfahani and Kuhn (2018) and the convergence to a global optimum is ensured by its convex structure.
@@
The training approach for physics-informed neural networks (PINNs) primarily involves a weighted multi-objective function designed to narrow the gap between labels and model predictions while penalizing physical constraint violations. This is often achieved by simultaneously minimizing the residuals of a partial differential equation (PDE) and its boundary conditions (Cuomo et al. 2022). Suppose all NN training weights are collected in $\bm{\Theta}$, a typical PINN training program takes the form:
@@
where $\mathcal{L_{F}}$ and $\mathcal{L_{B}}$ are the functions enforcing the differential equations and boundary conditions, respectively, of the governing physical system, $\mathcal{L}$ is the regression objective between labelled data and model predictions, and $\hat{\mathbf{Y}}=\Phi(\mathbf{X},\bm{\Theta})$ (Cuomo et al. 2022). The scalar weights $\lambda_{\mathcal{F}}$, $\lambda_{\mathcal{B}}$, and $\lambda_{\text{data}}$ must be tuned empirically.
We note that the constraints are softly enforced by penalizing their violation in the objective. It is similar to considering the Lagrangian relaxation of the constrained training problem. The loss terms related to the PDE are derived through the automatic differentiation of the neural network’s outputs (Baydin et al. 2017). This complex loss function is then iteratively optimized through backpropagation within the network. While Cuomo et al. (2022) note that boundary conditions can be strictly imposed during training by forcing a part of the network to satisfy these conditions, see Zhu et al. (2021), they are typically integrated as penalty terms within the loss function. By using the augmented Lagrangian in the objective function and increasing the penalty terms related to constraint violations iteratively, it is theoretically possible to converge to a PINN which imposes hard inequality constraints during training (Lu et al. 2021). Physics-informed neural networks are characterized by their extensive modelling capabilities, facilitated by automatic differentiation, which allows for the approximation of continuous PDEs across various neural network architectures, e.g., CNNs (Gao et al. 2021), long short-term memory networks (Zhang et al. 2020), Bayesian NNs (Yang et al. 2021), and generative adversarial networks (Yang et al. 2020). Although limited to convex constraints, in this work, we demonstrate that our training methodology can enforce hard constraints on the shallow network’s output during training without complexifying its architecture or objective function. These constraints may include, but are not limited to, boundary constraints, ramping constraints, discrete PDEs, and constraints derived from discrete-time dynamical systems.
@@
As stated previously, prediction algorithms used in large-scale VPP applications must respect some inherent constraints. They need to be computationally scalable, they must be able to model nonlinearities, and the sensibility of the application requires trustworthiness in their predictions, e.g., robustness, interpretability, and performance guarantees. Regarding the forecast of non-residential buildings’ energy consumption, because grid operators do not generally have access to information on building activity, occupancy, and control, the system is approached as a blackbox and data-driven approaches tend to be the default option (Amasyali and El-Gohary 2018). Authors have used support vector regression (SVR) (Dong et al. 2005) and other hybrid methods leveraging both classification and linear regression to satisfy the aforementioned constraints (Amasyali and El-Gohary 2018). These methods are linear by parts and can approximate nonlinear distributions adequately. Statistical methods such as autoregressive integrated moving average (ARIMA) (Newsham and Birt 2010) and Gaussian processes have also proved to be good contenders as they even offer empirical uncertainty bounds (Weng and Rajagopal 2015). Many authors have proposed deep-learning methods, yet these propositions do not have much deployment acceptability as they fail to meet the industrial constraints and bloat the requirements for a safe ML pipeline. In this work, we show that our proposition is interesting for virtual power plant applications where the forecast of many distributed assets is used for critical decision-making, as we satisfy all industrial constraints better than other comparable models.
@@
Distributionally robust learning, in general, aims at protecting machine learning models from out-of-distribution outliers in the training set. These outliers tend to be adversarial to the prediction quality of ML models in validation and testing sets and being able to ignore them should increase models’ out-of-sample performance. As such, to correctly benchmark distributionally robust machine learning models, one must be able to generate outliers that are indeed out-of-distribution with respect to the original distribution and be able to do it regardless of the form of the original distribution. In the literature, we found that methodologies to achieve these two objectives in regression tasks are limited. For example, Chen and Paschalidis (2018) create a training set by randomly drawing samples from two distributions, the inlying and outlying ones, under probability $p\in(0,1)$ and 1p1-p, respectively. They used Gaussian distributions for both. In Qi et al. (2022), the authors task their model with the forecast of a real-world dataset. To control the adversity of the setting, they discard the real labels and create a synthetic version of them by using the prediction of a linear regression trained on the original dataset. They then tune an additive noise over the labels to simulate different levels of outlier corruption. In Liu et al. (2022), they notably propose generating corrupted synthetic data by integrating attack mechanisms into the generation to bias their dataset, i.e., selection bias and anti-causal bias. While their methodology is interesting, they must rely on multiple Gaussian distributions to design their attacks as intended. By leveraging the sliced-Wasserstein distance (Bonnotte 2013), a tractable approximation of the Wasserstein distance, our methodology generates outliers, does not rely on any assumptions of the original distribution, making it quite flexible, and ensures that outliers do not create tangible patterns.
@@ 1.3 Contributions
We propose a new methodology to benchmark distributionally robust (and regularized) machine learning models in a controlled synthetic environment. We design a method to introduce real out-of-sample outliers under the sliced-Wasserstein distance (Bonnotte 2013), a computationally tractable approximation of the Wasserstein distance. We combine this method with noise to corrupt machine learning datasets generated with standard optimization benchmark functions typically reserved for non-convex solver benchmarking. We show that our WaDiRo-SCNN training performs well empirically in highly corrupted environments: it is robust and conservative while being able to model non-convex functions. Finally, we test our model on a real-world VPP application, viz., the hourly energy consumption prediction of non-residential buildings in Montréal, Canada. For this application, we enforce the physical constraint that energy consumption is non-negative. We observe that this constraint greatly reduces the number of constraint violations during testing while having no significant precision drop with the original non-physics-constrained WaDiRo-SCNN.
@@
The rest of the paper is organized as follows. In Section 2, we introduce Wasserstein distributionally robust optimization and its tractable formulation for convex problems (Section 2.1) as well as shallow convex neural networks (Section 2.2). In Section 3 we present the main results of this paper, namely, WaDiRo-SCNNs. In Sections 4, 5, and 6 we address the theoretical out-of-sample performance of our proposition, physics-constrained SCNNs, and SCNN post-training verification, respectively. Finally, in Section 7 we present a methodology to generate outliers and in Section 8, we present all numerical experiments.
@@ 2.1 Distributionally Robust Optimization
$\inf_{\bm{\beta}}\sup_{\text{Q}\in\Omega}\ \mathbb{E}^{\text{Q}}[h_{\bm{\beta}}(\mathbf{z})],$ (DRO)
where $\mathbf{z}\in\mathcal{Z}\subseteq\mathbb{R}^{d},d\in\mathbb{N}$, is a random vector, $\mathcal{Z}$ is the set of all possible values of $\mathbf{z}$, Q is a distribution from the ambiguity set $\Omega$ that characterizes the distribution of the random vector $\mathbf{z}$, $\bm{\beta}\in\mathbb{R}^{p}$ is the $p$-dimension decision vector, and $h_{\bm{\beta}}(\mathbf{z}):\mathcal{Z}\times\mathbb{R}^{p}\to\mathbb{R}$ is the objective function when applying $\bm{\beta}$ on a sample (Chen and Paschalidis 2020, Mohajerin Esfahani and Kuhn 2018).
We are interested in a data-driven branch of DRO which defines the ambiguity set as a Wasserstein ball, a ball which contains all distributions within a Wasserstein distance from a reference distribution (Yue et al. 2022), centred on an empirical distribution containing $N$ samples: $\hat{\text{P}}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{z}_{i}}$, where $\delta_{\mathbf{z}_{i}}$ is a Dirac delta function assigning a probability mass of one at each known sample $\mathbf{z}_{i}$. The empirical distribution is constructed with historical realizations of the random vector $\mathbf{z}$. Note that as the number of samples goes to infinity the empirical distribution weakly converges to the true distribution P (van der Vaart 2000, Chapter 19.2). The Wasserstein distance (WD), also called earth mover’s distance and closely related to the optimal transport plan (Kuhn et al. 2019), can be pictured as the minimal effort that it would take to displace a given pile of a weighted resource, e.g., dirt, to shape a second specific pile. The effort of a trip is given by the distance travelled and the weight of the carried resource. Specifically, it provides a sense of similarity between the two distributions. Formally, the probabilistic definition of the order-$t$ Wasserstein distance between probability distributions U and V on $\mathcal{Z}$ is:
$W_{\|\cdot\|,t}(\text{U},\text{V})=\inf_{\begin{subarray}{c}U\sim\text{U},\\ V\sim\text{V}\end{subarray}}\left(\mathbb{E}[\|U-V\|^{t}]\right)^{\frac{1}{t}},\quad t\geq 1,$
where the infimum is taken over all possible couplings of random vectors $U$ and $V$ which are marginally distributed as U and V, respectively (Panaretos and Zemel 2019). As such, the Wasserstein ball defines a space containing all possible distributions close to the distribution in its center. We refer interested readers to the open-access book of Panaretos and Zemel (2020) which covers the Wasserstein distance thoroughly while offering interesting statistical insights.
@@
$\inf_{\beta}\sup_{\text{Q}\in\Omega}\$$\mathbb{E}^{\text{Q}}[h_{\beta}(\mathbf{z})]$ (WDRO)
$\text{with}\ \Omega=$ $\{\text{Q}\in\mathcal{P}(\mathcal{Z}):W_{\|\cdot\|,t}(\text{Q},\hat{\text{P}}_{N})\leq\epsilon\},$
where $\Omega$ is the order-$t$ Wasserstein ball of radius $\epsilon>0$ centred on $\hat{\text{P}}_{N}$, and $\mathcal{P}(\mathcal{Z})$ is the set of all probability distributions with support on $\mathcal{Z}$. In a way, $\epsilon$ can be interpreted as a hyperparameter that controls our conservatism towards $\hat{\text{P}}_{N}$ because a greater radius admits distributions farther from $\hat{\text{P}}_{N}$ into the ambiguity set. This problem is not tractable (Mohajerin Esfahani and Kuhn 2018).
@@
$\inf_{\bm{\beta}}\sup_{\text{Q}\in\Omega}\ \mathbb{E}^{\text{Q}}[h_{\bm{\beta}}(\mathbf{z})]=$ $\inf_{\bm{\beta}}\kappa\epsilon+\frac{1}{N}\sum_{i=1}^{N}h_{\bm{\beta}}(\mathbf{z}_{i})$ (TWDRO)
$\,\text{s.t.}\ \kappa=\sup\{\|\bm{\theta}\|_{*}:h_{\bm{\beta}}^{*}(\bm{\theta})<+\infty\},$
where $\|\cdot\|_{*}$ is the dual norm, $\|\bm{\theta}\|_{*}\equiv\sup_{\|\mathbf{z}\|\leq 1}\bm{\theta}^{\top}\mathbf{z}$, and $h_{\bm{\beta}}^{*}(\bm{\theta})$ is the convex conjugate of $h_{\bm{\beta}}(\mathbf{z})$, $h_{\bm{\beta}}^{*}(\bm{\theta})=\sup_{\mathbf{z}}\{\bm{\theta}^{\top}\mathbf{z}-h_{\bm{\beta}}(\mathbf{z})\}$ (Boyd and Vandenberghe 2004).
Training machine learning models with (TWDRO) has been proved to be a generalization of regularization (Shafieezadeh-Abadeh et al. 2019). By training over the worst-case expected distribution in a Wasserstein ball centred on the training dataset, we can expect to avoid overfitting on the empirical distribution, thus yielding increased robustness against outliers, noise, and some adversarial attacks.